OpenStudy (anonymous):

What is the sum of 200+40+8+.....

OpenStudy (dean.shyy):

200/5 = 40 40/5 = 8 8/5 = ? So, the answer would be 200 + 40 + 8 + (8/5) = x Find x.

OpenStudy (anonymous):

249.6